OpenStudy (anonymous):

A survey was conducted in which 250 families were asked: How many bathrooms do you have in your household? The results are shown below. One Bathroom Two Bathrooms Three or More Bathrooms One bedroom 132 Two Bedroom 94 3 bedrooms 24 What is the probability that a randomly selected family has more than one bathroom? 0.528 0.004 0.472 0.376 0.096

OpenStudy (anonymous):

You've mixed up bathroom and bedroom in your question so I'm just going with bathroom for my answer. The probability of having more than one bathroom is 1-P(one bathroom) = 1-132/250=0.472

OpenStudy (anonymous):

yepp 0.472

OpenStudy (anonymous):

it one every family has 1

OpenStudy (anonymous):

its not asking about one bathroom, if how many have more then one, but thank you for your help